A manufacturing firm employs three analytical plan for the design and development of a particular product. For cost reasons, all three are used at varying times. In fact, plan 1, 2 and 3 are used for 30%, 25%, and 45% of the products, respectively. The defect rate is different for the three procedures as follows: P(D|P1) = 0.01, P(D|P2) = 0.03, P(D|P3) = 0.02
If a random product is observed and found to be defective,and is due to plan 1.
 

Grade:12th pass

1 Answers

Khimraj
5 years ago
possible outcome = 0.3*0.01 = 0.003
total outcome = 0.3*0.01 + 0.25*0.03 + 0.45*0.02 = 0.
so probability = 30/195 = 10/65  = 2/13
Hope it clears.